OpenStudy (anonymous):

Simplify ln(1/e) + lnAB This is a natural log problem right? Not really sure how to start it off though :/ But would the second part of the problem, lnAB, simplify to lnA + lnB ? :/

OpenStudy (dumbcow):

yes you are right about lnAB = lnA +lnB for first part, ln(1/e) = ln(e^-1) = -1*ln(e) = -1

OpenStudy (anonymous):

so this becomes: -1 + lnA + lnB ?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

okay yay :) so is that as simplified as it can get?

OpenStudy (dumbcow):

yeah well it depends what they mean by simplify, you can't break down the logs any further but you could combine everything into 1 log?? is this combining logs or expanding logs?

OpenStudy (anonymous):

not sure... it just says to simplify completely... :/

OpenStudy (anonymous):

so would that mean to break it up as much as possible?

OpenStudy (dumbcow):

yes i assume so

OpenStudy (anonymous):

okay great! thank you :)